OpenStudy (anonymous):

Write the equation of the line with a slope 3/7 passing through the point 9,-2.

OpenStudy (anonymous):

y=mx+b You have m(slope)= 3/7 And you have (x, y) (9,-2) x=9 and Y=-2 We need to find b. So we plug all this information in to the equation y=mx+b -2=(3/7)(9)+b Solve that and you will get b. Once you have b, plug it in to the equation, along with the slope.

OpenStudy (anonymous):

you wouldnt use y-y=m(x-x)??

OpenStudy (anonymous):

yes u can use that

OpenStudy (anonymous):

how would i set it up?

OpenStudy (anonymous):

u have ur gradient m = 3/7

OpenStudy (anonymous):

and ur point (9,-2) so let \[x _{1}=9 \] and \[y _{1}=-2\] and then plug ur values in the equation \[y=m(x-x _{1})+y _{1}\]

OpenStudy (anonymous):

dus that help?

OpenStudy (anonymous):

once i solve it, would that be my answer?

OpenStudy (anonymous):

yep.u have to get a linear equation

OpenStudy (anonymous):

ok thanks alot

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

is 2x-9y= -3/7 the correct answer?

OpenStudy (anonymous):

actually it is y=(3/7)x-(41/7) or 7y=3x-41

OpenStudy (anonymous):

check the equation i gave you and see if u have done the plugging and calculations accurately

OpenStudy (anonymous):

or rather show me step by step how u do it